Introduction to Poisson and Exponential distribution, Memoryless Property

You are working to statistically model the location of defects on the surface of a 3D printed material. After analyzing for one month you found out that the number of defects in the material follows a Poisson process, with an average of one defect is found every 10µm2 area. One of your lab mates is willing to validate the information. He selects 60 µm2 are on the surface to study the location of defects. Using this information provided, answer question 1.

1. What is the probability that your lab mate finds at least 4 defects on the surface area he is observing (round to 3 digits)?
Answer #1

lambda = 6

P(X >= 4)
= 1 - P(X <= 3)
= 1 - (P(X = 0) + P(X = 1) + P(X = 2) + P(X = 3) )
= 1 - (e^-6 + e^-6*6^1 + e^-6*6^2/2! + e^-6*6^3/3!)
= 0.8488

Ans: 0.849